# Universal Approximation Theory: Foundations for Parallelism in Neural Networks

### v1

@@ 5 Conclusion
In this paper, we propose a parallel network based on the Universal Approximation Theorem (UAT) and validate its effectiveness on large datasets. We analyze the characteristics of both serial and parallel networks. Serial networks, especially those based on residual connections, often have higher degrees of parameter freedom, leading to stronger generalization capabilities. However, as the network depth increases, the inference time also becomes longer. On the other hand, parallel networks trade computational resources for time, meaning their runtime is not affected by the number of network layers. This allows for the design of multi-layer parallel networks that can replace serial networks to achieve fast inference. Additionally, we analyze the impact and reasons data have on deep learning. From the UAT perspective, training on large datasets essentially enhances the model’s ability to fit data.
Appendix A The UAT Format of Transformer
A.1 The UAT Format of Transformer-based ViTs
Eq.(1) and (2) give an example of the matrix-vector format of MHA and FFN. Eq. (3) and Eq. (4) give the corresponding UAT format of $\mathbf{x}_{i+1}$ and $\mathbf{x}_{i+2}$. It is obvious that the bias and weight terms in the UAT corresponding to Transformed can change dynamically based on the input.
Due to the complexity of the variables involved, we will combine some terms to align the resulting equations with the standard UAT form. To distinguish these parameter variables, we will keep the original variables unchanged. If a parameter variable is influenced only by other variables, we will denote it with a bar on top, such as $\bar{\mathbf{W}}$. If it is influenced by the input $\mathbf{x}_{i}$, we will mark it with a hat, such as $\hat{\mathbf{W}}$.
$MHA(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}$ (1)
$FFN(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}\$ (2)
$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma[\mathbf{W}^{\prime}_{i,2}(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i})+\mathbf{b}^{\prime}_{i,2}]+\mathbf{b}^{\prime}_{i,3}$ (3)
$=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\underline{\mathbf{W}^{\prime}_{i,2}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}$
$=\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}$
$\mathbf{x}^{\prime}_{i+2}$ $=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i+1}+\mathbf{W}^{\prime}_{i+1,3}\sigma(\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{x}^{\prime}_{i+1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}$ (4)
$=\mathbf{W}^{\prime}_{i+1,1}[\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}]$
$+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,1}}[\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{b}^{\prime}_{i,3}]+\mathbf{b}^{\prime}_{i+1,2}\}+\mathbf{b}^{\prime}_{i+1,3}$
$=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,3}}$
$+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\underline{\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\underline{\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{W}^{\prime}_{i,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\underline{\hat{\mathbf{W}^{\prime}_{i+1,1}}\mathbf{b}^{\prime}_{i,3}+\mathbf{b}^{\prime}_{i+1,2}}\}+\mathbf{b}^{\prime}_{i+1,3}$
$=\hat{\mathbf{W}^{\prime}_{i+1,2}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{W}^{\prime}_{i+1,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,4}}\mathbf{x}^{\prime}_{i}+\underline{\hat{\mathbf{W}^{\prime}_{i+1,5}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\hat{\mathbf{b}^{\prime}_{i+1,2}}}\}+\underline{\hat{\mathbf{b}^{\prime}_{i,3}}+\mathbf{b}^{\prime}_{i+1,3}}$
$=\hat{\mathbf{W}^{\prime}_{i+1,2}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{W}^{\prime}_{i+1,3}}\sigma(\hat{\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i,2})+\mathbf{W}^{\prime}_{i+1,3}\sigma\{\hat{\mathbf{W}^{\prime}_{i+1,4}}\mathbf{x}^{\prime}_{i}+\hat{\mathbf{b}^{\prime}_{i+1,4}}\}+\hat{\mathbf{b}^{\prime}_{i+1,5}}$

### v4

@@ 5 Conclusion
In this paper, we propose a parallel network based on the UAT and validate its effectiveness on large datasets. We analyze the characteristics of both serial and parallel networks. Serial networks based on residual connections, often have higher degrees of parameter freedom, leading to stronger generalization capabilities. However, as the network depth increases, the inference time also becomes longer. On the other hand, parallel networks trade computational resources for time, meaning their runtime is not affected by the number of network layers. This allows for the design of multi-layer parallel networks that can replace serial networks to achieve fast inference. Additionally, we analyze the impact and reasons data have on deep learning. From the UAT perspective, training on large datasets essentially enhances the model’s ability to fit diverse data.
Appendix A The UAT Format of Transformer-based ViTs
In this section, we will demonstrate that the mathematical form of single-layer and two-layer Transformers is equivalent to the UAT. Firstly, the basic form of a Transformer is shown in Figure 2 of main text. We represent the MHA and FFN in the Transformer in matrix-vector form as Eq. (10) and Eq. (11), respectively (for specific details, refer to UAT2LLMs). To distinguish between the operations and variables in the Transformer and those in the matrix-vector form, we add a prime () symbol to the operations and variables in the matrix-vector form. It is important to emphasize that the parameters of the MHA are influenced by the input.
$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (10)
$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (11)
Next, the result of a single-layer Transformer $\mathbf{x}^{\prime}_{1}$ is:
$\mathbf{x}^{\prime}_{1}$ $=\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{W}^{\prime}_{1,3}\sigma[\mathbf{W}^{\prime}_{1,2}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0})+\mathbf{b}^{\prime}_{1,2}]+\mathbf{b}^{\prime}_{1,3}$ (12)
$=\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{W}^{\prime}_{1,3}\sigma(\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})+\mathbf{b}^{\prime}_{1,3}$
Let $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, so we obtain:
$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (13)
where $\mathbf{W}^{\prime}_{1,1}$ and $\mathbf{W}^{\prime}_{1,2}$ are parameters influenced by the input, thus having a degree of freedom of 2.
$\mathbf{x}^{\prime}_{2}$ $=\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})+\mathbf{b}^{\prime}_{2,3}$ (14)
$=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,3})+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})$
=$=\{\mathbf{W}^{\prime}_{2,1}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,3}\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,2}\}$
=$=\{\mathbf{W}^{\prime}_{2,1}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,3}+\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}({\mathbf{W}}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,2}+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
=$=\{(\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}})+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}\}$
In Eq. (14), we provide the matrix-vector format of 2-layer Transformer. Let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}$, $\mathbf{b}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$. Thus, Eq. (14) can be written as:
$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (15)
$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
where $\mathbf{W}^{\prime}_{2,1}$,$\mathbf{b}^{\prime}_{2,1}$,$\mathbf{W}^{\prime}_{1,2}$,$\mathbf{W}^{\prime}_{1,3}$,$\mathbf{W}^{\prime}_{2,3}$, and $\mathbf{b}^{\prime}_{2,2}$ are influenced by the input, so the degree of freedom is 6.
Another issue is that in $\mathbf{x}^{\prime}_{1}$ and $\mathbf{x}^{\prime}_{2}$, there is a term $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ and $(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})$ which doesn’t generally conform to the standard terms in UAT. We can consider this term as a special case and regard it as a layer in UAT. Firstly, when the number of network layers increases, it effectively increases $N$ in UAT. Given UAT’s strong approximation capabilities, adding multiple instances of this term can be viewed as a linear transformation, thus it doesn’t affect the overall approximation power of UAT. Furthermore, this term can be considered a special form in UAT. Taking ReLU as an example activation function, $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ can be expressed as $\text{ReLU}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})-\text{ReLU}(-(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3}))$.
In summary, we have shown that the mathematical forms of one-layer and two-layer Transformers can be seen as 2 and three-layer UAT.
Appendix B The Degree of Freedom of Different Layers Transformer
According to UAT2LLMs, the UAT format of multi-layer Transformer can be written as:
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (16)
where $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$. The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by a $j$-layer of UAT with $\mathbf{x}_{0}$ as input. All $\mathbf{W}^{\prime}_{i,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input. Because they are affected by the weight of MHA in each layer. Why does the MHA operation affect the weight parameters in a linear manner? According to Eq. (12) and Eq. (13), this is because $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ represents the MHA parameters, which are influenced by the input, while $\mathbf{W}^{\prime}_{1,2}$ is a defined parameter. As a result, the parameter $\mathbf{W}^{\prime}_{1,2}$ in Eq. (13) becomes input-dependent, and this influence occurs through a linear transformation. Similarly, in a multi-layer Transformer, multiple layers of weight parameters are derived through similar transformations.
Therefore, it is easy to deduce that the degrees of freedom for Transformers with 1, 2, 3, and 6 layers are 2, 6, 9, and 18, respectively. For Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24, the degrees of freedom are shown in Table 1. At first glance, it seems that the degrees of freedom for Para-F-1-24, Para-F-2-24, Para-F-3-24, and Para-F-6-24 all exceed that of Para-F-6-6. Why then does only Para-F-6-24 consistently outperform Para-F-6-6?
Para-F-1-24 Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
Degree of Freedom 2*24 18*1  6*12 9*8 18*4
Table 1: The degree of freedom of Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, Para-F-6-24. We use the the degree of freedom of serial Transformer with the depth 1, 6, 2, and 3 to multiply the number of serial Transformer.
The reason lies in the nature of the degrees of freedom. Taking $\mathbf{W}^{\prime}_{2,2}$ and $\mathbf{b}^{\prime}_{2,2}$ from Eq. (15) as examples, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ is an input-dependent parameter from MHA, and $\mathbf{W}^{\prime}_{2,2}$ is a defined parameter, becoming fixed post-training. Thus, the degrees of freedom for $\mathbf{W}^{\prime}_{2,2}$ in Eq. (15) are imparted through a linear transformation with the MHA parameter $\mathbf{W}^{\prime}_{1,1}$. Additionally, $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$. Hence, $\mathbf{b}^{\prime}_{2,2}$ in Eq. (15) can be understood as being approximated by a two-layer UAT, showcasing stronger dynamic transformation capabilities based on inputs. Due to UAT’s powerful capacity, $\mathbf{b}^{\prime}_{2,2}$ can dynamically fit specific functions based on inputs.
In a single-layer Transformer, as seen in Eq. (12), $\mathbf{b}^{\prime}_{1,3}$ and $\mathbf{b}^{\prime}_{1,2}$ are defined parameters, so the number of bias terms approximated using UAT is zero. For a two-layer Transformer, as shown in Eq. (15), the number of bias terms approximated using UAT is two, specifically $\mathbf{b}^{\prime}_{2,1}$ and $\mathbf{b}^{\prime}_{2,2}$, with the UAT layers approximating these bias terms being 1 and 2, respectively. According to Eq.(16), in a Transformer with $i+1$ layers, $\mathbf{b}_{i+1,1}$ is approximated by a single-layer UAT, while $\mathbf{b}^{\prime}_{j,2}$ is approximated by a UAT of $j$ layers for $j>1$. It is easy to deduce that for a three-layer Transformer, the number of bias terms approximated using UAT is three, with the UAT layers being 1, 2, and 3. For a six-layer Transformer, the number of bias terms approximated using UAT is six, with the UAT layers being 1, 2, 3, 4, 5, and 6. The UAT layers of bias terms for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are shown in Table 2.
Para-F-1-24 Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
UAT Layers (0)*24 (1+2+3+4+5+6)*1 (1+2)*12 (1+2+3)*8 (1+2+3+4+5+6)*4
Table 2: The UAT layer count corresponding to the bias terms in Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 is calculated by summing the UAT layers associated with each bias term within a serial Transformer. Since each serial Transformer has multiple bias terms, with each term corresponding to different UAT layer counts, we represent the UAT layers by summing the UAT layers for each bias term. For example, ”1+2+3” indicates that three bias terms are approximated using UAT, with 1, 2, and 3 layers respectively. Additionally, since multiple serial Transformers are used in parallel, we multiply the sum by the number of parallel Transformers.
In summary, more layers mean more degrees of freedom in bias terms, leading to stronger model fitting capabilities. To achieve performance comparable to that of deep serial networks while ensuring model fitting ability and fast inference, parallel networks need to strike a balance between network depth and layer count. Parallel networks are designed for extremely large models. For instance, if a serial network with 1 million layers can be replaced by a parallel network with a depth of 10 and 100 million layers, the speedup factor can reach up to 100,000 times. Sure, we could improve the design of the Para-former to achieve speedup computing with fewer layers.